OpenStudy (anonymous):

Sharon has 8 cups of granola. She eats 2/3 cup each day. How long will her granola last? A. 5 1/3 days B. 8 2/3 days C. 12 days D. 18 days

OpenStudy (anonymous):

I think B?

OpenStudy (anonymous):

Type your reply

OpenStudy (wilsondanielle):

try finding and equation that will relate the cups of granola to what she eats a day. For example, the first day she has consumed 2/3 of a cup. The second, she has consumed a total of 4/3 (or 1 cup and a third) of granola a day. Once you find the relationship, you can solve for y (total granola consumed) = 8 cups

OpenStudy (anonymous):

so C?

OpenStudy (anonymous):

brb im goiing to eat

OpenStudy (wilsondanielle):

Yes. I got 2x/3 = y so 2x/3 = 8 2x = 8*3 2x = 24 x=24/2 x=12

OpenStudy (anonymous):

cool thank you